OpenStudy (anonymous):

Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation 1 4-3x=7x-6 Equation 2 8+4x=-10+x The solution to equation #1 is larger than the solution to equation #2. The solution to equation #1 is smaller than the solution to equation #2. The solution to equation #1 is the same as the solution to equation #2. None of the statements above describe solutions to equations shown. The solution to equation #1 is the same as the solution to equation #2. The solution to equation #1 is smaller than the solution to equation #2. None of the statements above describes solutions to equations shown. The solution to equation #1 is larger than the solution to equation #2.

OpenStudy (anonymous):

The solution to equation #1 is larger than the solution to equation #2.

OpenStudy (anonymous):

how

OpenStudy (anonymous):

the ans to first equation is x=1, for the other it is -38/3

OpenStudy (anonymous):

4-3x=7x-6 10x=10 x=1 28+4x=-10+x 3x=-38 x=-38/3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome